# Mixed-Precision Embeddings for Large-Scale Recommendation Models

### v1

@@ 1.1. Background
In the deep learning era, embedding techniques have become essential components of large databases due to their effectiveness and efficiency in representing and retrieving information (Günther, 2018; Lubarsky et al., 2023). By encoding discrete entities, such as words, items, or graph nodes, into continuous vector spaces, embeddings facilitate more efficient storage, retrieval, and processing within large databases. These techniques are now widely applied across various domains, including natural language processing (Mikolov et al., 2013; Patil et al., 2023), graph learning (Grover and Leskovec, 2016; Li et al., 2024b), and recommender systems (Li et al., 2024a; Zhang et al., 2021; Yang et al., 2024). In particular, embedding techniques have a great impact on recommender systems, where the number of required embeddings far exceeds that of other domains. Specifically, the input data in recommender systems typically consists of various attributes, such as user_id, age, and occupation, each with multiple possible values. These attributes are often referred to as feature fields, while their specific values are known as categorical features, or simply features. Each categorical feature will be represented by a unique embedding vector, which enhances the modeling of similarities and interactions among features, thereby improving the accuracy of recommendations. However, there are usually a vast number of categorical features in recommender systems. For example, an ordinary advertising system in Baidu contains billions of features (Xu et al., 2021), whereas the vocabulary size of the large language model LLaMA 3 (Dubey et al., 2024) is only 128,000, with the former being several orders of magnitude larger.
@@
The typical architecture of deep learning recommendation models (DLRMs) includes an embedding table that converts categorical features into dense vector representations, followed by a feature interaction network that processes these embeddings to generate predictions, as illustrated in Figure 1(a). Numerous studies have refined the structure of the feature interaction network to improve prediction accuracy (Guo et al., 2017; Wang et al., 2017). Despite these improvements, the interaction networks often have relatively shallow layers and a limited number of parameters, while the majority of model parameters are concentrated in the embedding table. Especially when handling billions of categorical features, the storage of the embedding table can even reach 10 TB (Xu et al., 2021). Therefore, developing techniques to compress embeddings while maintaining model accuracy have become a crucial issue for recommender systems.
@@
Regarding model compression, quantization is undoubtedly one of the most effective technique (Nagel et al., 2021), which reduces memory usage by representing parameters with fewer bits. There are two primary methods of quantization: post-training quantization (PTQ) (Guan et al., 2019; Nagel et al., 2020) and quantization-aware training (QAT) (Esser et al., 2020; Bhalgat et al., 2020; Nagel et al., 2022). PTQ applies quantization after the training process, while QAT incorporates a quantization operator during training. As shown in Figure 1(b), QAT inserts a fake quantization operator into the forward propagation, allowing the model to better adapt to quantization and thus achieving higher accuracy than PTQ. During backpropagation, the learnable parameters are updated using the Straight-Through Estimator (STE) (Hinton, 2012), which approximates the gradient of the non-differentiable quantization operator as 1, thereby enabling end-to-end optimization. After training, the parameters will be quantized and then stored in a low-precision format for deployment.
@@ 1.2. Motivation and Contributions
An embedding table can be represented by a matrix $\boldsymbol{E}\in\mathbb{R}^{n\times d}$, where $n$ is the number of categorical features, and $d$ is the dimension of each embedding vector. As discussed by Zhang et al. (2024), existing methods generally compress either the rows or columns of the embedding table. Row compression methods involve techniques like feature selection (Lyu et al., 2023; Lai et al., 2023; Guo et al., 2022; Jia et al., 2024) and hashing (Shi et al., 2020; Zhang et al., 2020; Yan et al., 2021; Zhang et al., 2024). For example, OptFS (Lyu et al., 2023) filters out unimportant features to reduce the number of rows in the embedding table, while Zhang et al. (2020) propose sharing embeddings among infrequent features through hash functions, thereby reusing rows. On the other hand, dimension compression methods involve techniques like mixed-dimension embeddings (Sun et al., 2020; Zhao et al., 2021a, b; Lyu et al., 2022) and pruning (Liu et al., 2021; Deng et al., 2021; Xiao et al., 2022). For example, CpRec (Sun et al., 2020) assigns varying dimensions to embeddings of different features, while PEP (Liu et al., 2021) prunes redundant parameters within each feature embedding by learning a threshold. A key characteristic of these methods is their ability to explicitly or implicitly assess feature importance and allocate storage space accordingly.
@@
In addition to row and column perspectives, parameter precision is another critical perspective of embedding compression, closely related to quantization techniques. Guan et al. (2019) were the first to apply PTQ to embedding tables. Later studies (Xu et al., 2021; Yang et al., 2020; Li et al., 2023) have focused on the low-precision training (LPT) paradigm. Unlike QAT, which retains full-precision parameters during training, LPT keeps embeddings in a low-precision format throughout training to reduce training memory usage. However, at lower precision levels, LPT becomes impractical due to significant accuracy loss. For example, ALPT (Li et al., 2023) achieves lossless compression only at 8-bit precision. Note that recommender systems typically cannot tolerate any degradation in accuracy. Beyond these efforts, there has been little in-depth research on quantization, particularly QAT, within recommender systems. This scarcity is largely attributable to the unique characteristic of embedding tables in recommender systems, which render quantization techniques less prevalent compared to domains such as computer vision (Esser et al., 2020) and natural language processing (Hou et al., 2019). Specifically, embedding tables often contain a vast number of features with varying levels of importance. Yet, quantization typically applies uniform precision across all embeddings, failing to distinguish feature importance. As a result, numerous less important features are assigned the same precision as more significant ones, even though they could be represented with lower precision, leading to inefficient compression.
@@
In this paper, we aim to improve the quantization of embedding tables by incorporating the ability to distinguish feature importance. Previous studies have linked feature importance to the embedding dimension (Sun et al., 2020), the utilization of the feature (Lyu et al., 2023) and so on. Similarly, we propose to associate feature importance with embedding precision. More precisely, our objective is to find the most appropriate precision for each feature embedding to balance model accuracy with memory efficiency. It is worth noting that feature selection can be viewed as a binary case of mixed-precision embeddings, where the search space includes only zero-precision and full-precision options. Therefore, we believe mixed-precision embeddings can achieve finer-grained compression than feature selection methods.
@@
To solve the problem of searching embedding precision, we propose a novel algrithom, termed Mixed-Precision Embeddings (MPE). There are two main challenges in building MPE. The first challenge is the vast number of categorical features. DLRMs typically involve millions of categorical features, resulting in an excessively large search space when searching precision for each feature individually. To mitigate this issue, we propose grouping features and optimizing precision for each group, which substantially reduces the search space complexity. Previous research has identified feature frequency as a reliable measure of feature importance (Zhang et al., 2020; Lai et al., 2023). Therefore, we suggest grouping features by frequency to ensure that features within the same group have similar levels of importance, thereby minimizing conflicts from shared precision. The second challenge is to determine the precision (i.e., quantization bit-width) for each feature group. However, precision search is inherently a discrete problem and cannot be effectively solved by traditional gradient descent algorithms. To achieve end-to-end optimization, we propose learning a probability distribution over candidate bit-widths for each group, where output embeddings will be represented as the expected values of quantized embeddings across all candidate bit-widths. Additionally, we apply regularization to the expected bit-width of groups to minimize memory usage. By balancing the objective function with this regularization, we can effectively learn the optimal probability distribution over candidate bit-widths. Finally, we sample the precision based on the optimized probability distribution and then retrain the model to enhance accuracy.
@@ 3. Mixed-Precision Embeddings
In this section, we first provide a formal definition of embedding precision search problem in Section 3.1. Then, we introduce how to reduce the complexity of the search space in Section 3.2 and how to search precision for feature embeddings in Section 3.3. Finally, we illustrate how to generate mixed-precision embeddings with the searched precision in Section 3.4.
@@ 3.1. Problem Formulation
It is a consensus that categorical features in recommender systems have varying levels of importance (Zhang et al., 2020; Lyu et al., 2023). Previous studies leveraged this variation in feature importance to reduce the parameters associated with less important features (Zhang et al., 2020; Zhao et al., 2021a; Lyu et al., 2023; Lai et al., 2023). For example, Sun et al. (2020) reduces the embedding dimension of less important features through heuristic rules, while AutoEmb (Zhao et al., 2021a) automatically searches embedding dimensions with reinforcement learning algorithms. Additionally, AdaEmbed (Lai et al., 2023) uses gradient information and feature frequencies to access feature importance and then prunes unimportant features. Despite these advancements, no prior work has systematically combined feature importance with corresponding embedding precision. In this paper, to achieve more efficient embedding compression, we propose searching an optimal precision level for each feature, which refers to the smallest bit-width without losing prediction accuracy.
Given a set of candidate precision levels (i.e., bit-widths), $\mathbb{B}=\{b_{1},b_{2},...,b_{m}\}$, we aim to determine $\boldsymbol{B}\in\mathbb{B}^{n\times 1}$ for the embedding table $\boldsymbol{E}\in\mathbb{R}^{n\times d}$, where each element of $\boldsymbol{B}$ represents the precision of the corresponding embedding in $\boldsymbol{E}$. Specifically, the problem of mixed-precision embeddings can be formally defined as follows:
@@ 3.3. Learnable Probability Distribution
In the previous section, we reduced the search space by grouping features. The remaining challenge is determining the optimal precision for each feature group. However, optimizing the embedding precision is more complex than solving discrete problems like feature selection or embedding dimension search. Feature selection is a simple decision about whether to include a feature, while embedding dimension search involves selecting from multiple candidate dimensions. In contrast, mixed-precision embeddings presents two discrete optimization problems: selecting from multiple candidate bit-widths and managing non-differentiable quantization operators.
Inspired by traditional QAT algorithms, we employ the STE to handle the non-differentiability of quantization operators. Further, to optimize the precision of different feature groups in an end-to-end manner, we convert the one-hot selection of candidate bit-widths into the learning of a probability distribution. Specifically, for a given set of candidate bit-widths, $\mathbb{B}=\{b_{1},b_{2},...,b_{m}\}$, we aim to learn a probability distribution $\boldsymbol{p}=\{p_{1},p_{2},...,p_{m}\}$ for each feature group, where $p_{i}$ denotes the likelihood of selecting $b_{i}$ as the bitwidth. To ensure the probabilities sum to 1, we maintain a learnable vector $\gamma=\{\gamma_{1},\gamma_{2},...,\gamma_{m}\}$ for each group and calculate the probabilities using the softmax function, as shown below:
@@
Using the probability distribution above, each feature in a group has a specific probability of being quantized at various bit-widths. Therefore, we use the expected outcomes of quantizing the feature embedding at different bit-widths as the final quantized embedding, as shown Figure 2. Specifically, for an embedding vector $\boldsymbol{e}$, the final quantized embedding $\bar{\boldsymbol{e}}$ will be computed as follows:
@@ 3.4. Precision Sampling and Retraining
Through the optimization process illustrated in Figure 2, we can derive the probability distribution over candidate bit-widths, which allows us to determine the final bit-width for each group. However, the bit-width with the highest probability may not necessarily be the most appropriate choice. Specifically, while a higher bit-width may not correspond to the highest probability, it can still contribute a considerable portion of high-precision representation to the final embedding. In this context, this feature group should be recognized as one that necessitates a higher bit-width. To this end, we select the highest bit-width with a probability exceeding the threshold of $\frac{1}{2m}$ as the final precision for each group, as outlined below:
@@
On the other hand, directly quantizing the embedding parameters using the sampled bit-width may create discrepancies between the quantized embeddings and those generated during training, thus affecting the model accuracy. This discrepancy arises because the embeddings used during training result from combining quantization outcomes across all candidate bit-widths. To mitigate accuracy loss, we implement a retraining process utilizing the selected bit-widths to train mixed-precision embeddings from scratch. Following the Lottery Ticket Hypothesis (LTH) (Frankle and Carbin, 2019), previous studies (Liu et al., 2021; Zhao et al., 2021b) will reset the model parameters to the initialization values for retraining. However, to leverage the information from the initial search phase, we employ the optimized step size, offset, and the feature interaction network to initialize the model in the retraining phase, aiming to enhance training outcomes. Note that only the embedding parameters are set back to the initialized values of the search phase during retraining.
@@ 4. Implementation
During inference, MPE only require storing embeddings in mixed-precision formats and do not involve mixed-precision computations. Once the low-precision embeddings are retrieved, they must be dequantized into full-precision parameters before being fed into subsequent networks. Note that the input to the embedding table is highly sparse, and the embeddings retrieved during forward propagation constitute only a small proportion of the total. As a result, the storage overhead for dequantized embedding parameters is negligible. However, since frameworks such as PyTorch do not natively support data formats such as Int-2 and Int-3, we concatenate each embedding vector at the bit level and then divide them into the Int-16 format for storage. Upon feature retrieval, the corresponding Int-16 data will be converted into full-precision embeddings.

### v2

@@ 1. Introduction
[…]
The typical architecture of deep learning recommendation models (DLRMs) and the schematic diagram of quantization-aware training (QAT).
@@ 1.1. Background
In the deep learning era, embedding techniques have become essential components of large databases due to their effectiveness and efficiency in representing and retrieving information (Günther, 2018; Lubarsky et al., 2023). By encoding discrete entities, such as words, items, or graph nodes, into continuous vector spaces, embeddings facilitate more efficient storage, retrieval, and processing within large databases. These techniques are now widely applied across various domains, including natural language processing (Mikolov et al., 2013; Patil et al., 2023), graph learning (Grover and Leskovec, 2016; Li et al., 2024b), and recommender systems (Li et al., 2024a; Zhang et al., 2021; Yang et al., 2024). In particular, embedding techniques have a great impact on recommender systems, where the number of required embeddings far exceeds that of other domains. Specifically, the input data in recommender systems typically consists of various attributes, such as user_id, age, and occupation, each with multiple possible values. These attributes are often referred to as feature fields, while their specific values are known as categorical features or simply features. Each categorical feature will be represented by a unique embedding vector, which enhances the modeling of similarities and interactions among features, thereby improving the accuracy of recommendations. However, the amount of categorical features in recommender systems is significantly larger than that of discrete entities in other domains. For example, an ordinary advertising system in Baidu contains billions of features (Xu et al., 2021). In contrast, the vocabulary size of the large language model LLaMA 3 (Dubey et al., 2024) is only 128,000, with the former being several orders of magnitude larger. Therefore, embedding techniques are particularly important in recommender systems.
@@
The embedding techniques are commonly adopted as part of the deep learning recommendation models (DLRMs). The typical architecture of DLRMs includes an embedding table that converts categorical features into dense vector representations, followed by a feature interaction network that processes these embeddings to generate predictions, as illustrated in Figure 1(a). Numerous studies have refined the structure of the feature interaction network to improve prediction accuracy (Guo et al., 2017; Wang et al., 2017; Qu et al., 2016; Wang et al., 2021a; Lyu et al., 2022a; Xiao et al., 2017). Despite these improvements, the interaction networks often have relatively shallow layers and a limited number of parameters, while the majority of model parameters are concentrated in the embedding table. Especially when handling billions of categorical features, the storage of the embedding table can even reach 10 TB (Xu et al., 2021). Therefore, developing techniques to compress embeddings while maintaining model accuracy has become a crucial issue in deploying DLRMs.
@@
Regarding model compression, quantization is undoubtedly one of the most effective techniques (Nagel et al., 2021), which reduces memory usage by representing parameters with fewer bits. There are two primary methods of quantization: post-training quantization (PTQ) (Guan et al., 2019; Nagel et al., 2020) and quantization-aware training (QAT) (Esser et al., 2020; Bhalgat et al., 2020; Nagel et al., 2022). PTQ applies quantization after the training process, while QAT incorporates a quantization operator during training. As shown in Figure 1(b), QAT inserts a fake quantization operator into the forward propagation, allowing the model to better adapt to quantization and thus achieve higher accuracy than PTQ. During backpropagation, the learnable parameters are updated using the Straight-Through Estimator (STE) (Hinton, 2012), approximating the gradient of the non-differentiable quantization operator as 1, thereby enabling end-to-end optimization. After training, the parameters will be quantized and then stored in a low-precision format for deployment.
@@ 1.2. Motivation and Contributions
An embedding table can be represented by a matrix $\boldsymbol{E}\in\mathbb{R}^{n\times d}$, where $n$ is the number of categorical features, and $d$ is the dimension of each embedding vector. As discussed by Zhang et al. (2024), existing methods generally compress either the rows or columns of the embedding table. Row compression methods involve techniques like feature selection (Lyu et al., 2023; Lai et al., 2023; Guo et al., 2022; Jia et al., 2024) and hashing (Shi et al., 2020; Zhang et al., 2020; Yan et al., 2021; Zhang et al., 2024). For example, OptFS (Lyu et al., 2023) filters out unimportant features to reduce the number of rows in the embedding table, while Zhang et al. (2020) propose sharing embeddings among infrequent features through hash functions, thereby reusing rows. On the other hand, dimension compression methods involve techniques like mixed-dimension embeddings (Sun et al., 2020; Zhao et al., 2021a, b; Lyu et al., 2022b) and pruning (Liu et al., 2021; Deng et al., 2021; Xiao et al., 2022). For example, CpRec (Sun et al., 2020) assigns varying dimensions to embeddings of different features, while PEP (Liu et al., 2021) prunes redundant parameters within each feature embedding by learning a threshold. A key characteristic of these methods is their ability to explicitly or implicitly assess feature importance and allocate storage space accordingly.
@@
In addition to row and column perspectives, parameter precision is another critical perspective of embedding compression, closely related to quantization techniques. Guan et al. (2019) were the first to apply PTQ to embedding tables. Later studies (Xu et al., 2021; Yang et al., 2020; Li et al., 2023) have focused on the low-precision training (LPT) paradigm. Unlike QAT, which retains full-precision parameters during training, LPT keeps embeddings in a low-precision format throughout training to reduce training memory usage. However, at lower precision levels, LPT becomes impractical due to significant accuracy loss. For example, ALPT (Li et al., 2023) achieves lossless compression only at 8-bit precision. Note that recommender systems typically cannot tolerate any degradation in accuracy. Beyond these efforts, there has been little in-depth research on quantization, particularly QAT, within recommender systems. This scarcity is largely attributable to the unique characteristic of embedding tables in recommender systems, which render quantization techniques less prevalent compared to domains such as computer vision (Esser et al., 2020) and natural language processing (Hou et al., 2019). Specifically, embedding tables often contain many features with varying levels of importance. Yet, quantization typically applies uniform precision across all embeddings, failing to distinguish feature importance. As a result, numerous less important features are assigned the same precision as more significant ones, even though they could be represented with lower precision, leading to inefficient compression.
@@
In this paper, we aim to improve the quantization of embedding tables by incorporating the ability to distinguish feature importance. Previous studies have linked feature importance to the embedding dimension (Sun et al., 2020), the utilization of the feature (Lyu et al., 2023) and so on. Similarly, we propose to associate feature importance with embedding precision. More precisely, our objective is to find the most appropriate precision for each feature embedding to balance model accuracy with memory efficiency. It is worth noting that feature selection can be viewed as a binary case of mixed-precision embeddings, where the search space includes only zero-precision and full-precision options. Therefore, mixed-precision embeddings can achieve finer-grained compression than feature selection methods.
@@
To solve the problem of searching embedding precision, we propose a novel algorithm named Mixed-Precision Embeddings (MPE). There are two main challenges in building MPE. The first challenge is the vast number of categorical features. DLRMs typically involve millions of categorical features, resulting in an excessively large search space when searching precision for each feature individually. To mitigate this issue, we propose grouping features and optimizing precision for each group, substantially reducing search space complexity. Previous research has identified feature frequency as a reliable measure of feature importance (Zhang et al., 2020; Lai et al., 2023). Therefore, we suggest grouping features by frequency to ensure that features within the same group have similar levels of importance, thereby minimizing conflicts from shared precision. The second challenge is determining each feature group’s precision (i.e., quantization bit-width). However, precision search is inherently a discrete problem that traditional gradient descent algorithms cannot solve effectively. To achieve end-to-end optimization, we propose learning a probability distribution over candidate bit-widths for each group, where output embeddings will be represented as the expected values of quantized embeddings across all candidate bit-widths. Additionally, we apply regularization to the expected bit-width of groups to minimize memory usage. By balancing the objective function with this regularization, we can effectively learn the optimal probability distribution over candidate bit-widths. Finally, we sample the precision based on the optimized probability distribution and then retrain the model to enhance accuracy.
@@ 3. Mixed-Precision Embeddings
In this section, we first provide a formal definition of embedding precision search problem in Section 3.1. Then, we introduce how to reduce the complexity of the search space in Section 3.2 and how to search precision for feature embedding in Section 3.3. Finally, we illustrate how to generate mixed-precision embedding with the searched precision in Section 3.4.
@@
[…]
Learning process of the probability distribution over candidate bit-widths in MPE.
@@ 3.1. Problem Formulation
It is a consensus that categorical features in recommender systems have varying levels of importance (Zhang et al., 2020; Lyu et al., 2023). Previous studies leveraged this variation in feature importance to reduce the parameters associated with less important features (Zhang et al., 2020; Zhao et al., 2021a; Lyu et al., 2023; Lai et al., 2023). For example, Sun et al. (2020) reduces the embedding dimension of less important features through heuristic rules, while AutoEmb (Zhao et al., 2021a) automatically searches embedding dimensions with reinforcement learning algorithms. Additionally, AdaEmbed (Lai et al., 2023) uses gradient information and feature frequencies to access feature importance and then prunes unimportant features. Despite these advancements, no prior work has systematically combined feature importance with corresponding embedding precision. In this paper, to achieve more efficient embedding compression, we propose searching for an optimal precision level for each feature, which refers to the smallest bit-width without losing prediction accuracy.
Given a set of candidate precision levels (i.e., bit-widths), $\mathbb{B}=\{b_{1},b_{2},...,b_{m}\}$, we aim to determine $\boldsymbol{B}\in\mathbb{B}^{n\times 1}$ for the embedding table $\boldsymbol{E}\in\mathbb{R}^{n\times d}$, where each element of $\boldsymbol{B}$ represents the precision of the corresponding embedding in $\boldsymbol{E}$. Specifically, the problem of mixed-precision embedding can be formally defined as follows:
@@ 3.3. Learnable Probability Distribution
In the previous section, we reduced the search space by grouping features. The remaining challenge is determining the optimal precision for each feature group. However, optimizing the embedding precision is more complex than solving discrete problems like feature selection or embedding dimension search. Feature selection is a simple decision about whether to include a feature while embedding dimension search involves selecting from multiple candidate dimensions. In contrast, mixed-precision embedding presents two discrete optimization problems: selecting from multiple candidate bit-widths and managing non-differentiable quantization operators.
Inspired by traditional QAT algorithms, we employ the STE to handle the non-differentiability of quantization operators. Further, to optimize the precision of different feature groups in an end-to-end manner, we convert the one-hot selection of candidate bit-widths into the learning of a probability distribution. Specifically, for a given set of candidate bit-widths, $\mathbb{B}=\{b_{1},b_{2},...,b_{m}\}$, we aim to learn a probability distribution $\boldsymbol{p}=\{p_{1},p_{2},...,p_{m}\}$ for each feature group, where $p_{i}$ denotes the likelihood of selecting $b_{i}$ as the bit-width. To ensure the probabilities sum to 1, we maintain a learnable vector $\gamma=\{\gamma_{1},\gamma_{2},...,\gamma_{m}\}$ for each group and calculate the probabilities using the softmax function, as shown below:
@@
Using the probability distribution above, each feature in a group has a specific probability of being quantized at various bit-widths. Therefore, we use the expected outcomes of quantizing the feature embedding at different bit-widths as the final quantized embedding, as shown in Figure 2. Specifically, for an embedding vector $\boldsymbol{e}$, the final quantized embedding $\bar{\boldsymbol{e}}$ will be computed as follows:
@@ 3.4. Precision Sampling and Retraining
Through the optimization process illustrated in Figure 2, we can derive the probability distribution over candidate bit-widths, allowing us to determine each group’s final bit-width. However, the bit-width with the highest probability may not necessarily be the most appropriate choice. Specifically, while a higher bit-width may not correspond to the highest probability, it can still contribute a considerable portion of high-precision representation to the final embedding. In this context, this feature group should be recognized as one that necessitates a higher bit-width. To this end, we select the highest bit-width with a probability exceeding the threshold of $\frac{1}{2m}$ as the final precision for each group, as outlined below:
@@
On the other hand, directly quantizing the embedding parameters using the sampled bit-width may create discrepancies between the quantized embedding and that generated during training, thus affecting the model’s accuracy. This discrepancy arises because the embedding used during training result from combining quantization outcomes across all candidate bit-widths. To mitigate accuracy loss, we implement a retraining process using the sampled bit-widths to train mixed-precision embedding from scratch. Following the Lottery Ticket Hypothesis (LTH) (Frankle and Carbin, 2019), previous studies (Liu et al., 2021; Zhao et al., 2021b) will reset the model parameters to the initialization values for retraining. However, to leverage the information from the search phase, we employ the optimized step size, offset, and feature interaction network to initialize the model in the retraining phase, aiming to enhance training outcomes. Note that only the embeddings are set back to the initialized values of the search phase during retraining.
@@ 4. Implementation
During inference, MPE only requires storing embedding in mixed-precision formats and does not involve mixed-precision computations. Once the low-precision embeddings are retrieved, they must be dequantized into full-precision parameters before being fed into subsequent networks. Note that the input to the embedding table is highly sparse, and the embedding retrieved during forward propagation constitutes only a small proportion of the total. As a result, the storage overhead for dequantized embedding parameters is negligible. However, since frameworks such as PyTorch do not natively support data formats such as Int-2 and Int-3, we concatenate each embedding vector at the bit level and then divide them into the Int-16 format for storage. Upon feature retrieval, the corresponding Int-16 data will be converted into full-precision embedding.